OpenStudy (anonymous):

What is the 6th term in this geometric sequence? 480, 240, 120, 60, ... A. 15 B. 30 C. 7.5

sam (.sam.):

r=1/2 a=480 Tn=ar^(n-1) T6=(480)(1/2)^(6-1)=15